OpenStudy (nuccioreggie):

The graph shows the amount of money paid when purchasing bags of caramel corn at the zoo: A graph is shown. The x-axis is labeled from 0 to 9. The y-axis is labeled from 0 to 10. The four points are shown on the graph are on ordered pairs 0, 0 and 1, 3 and 2, 6 and 3, 9. These points are joined by a line. The label on the x-axis is Bags of caramel corn. The title on the y-axis is Total cost in dollars. Write an equation to represent the relationship between the total cost (y) and the number of bags of caramel corn (x). y = one-half x y = 2x y = 3x y = one-third x

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

looks like you multiply the first number by 3 to get the second number right?

OpenStudy (nuccioreggie):

ok i think the answer is B

OpenStudy (misty1212):

\[1\times 3=3\\ 2\times 3=6\\ 3\times 3=9\] i.e. you multiply each number by three

OpenStudy (misty1212):

interesting... why do you think you multiply each number by two?

OpenStudy (nuccioreggie):

o sorry is it A

OpenStudy (misty1212):

oh so you divide each number by two?

OpenStudy (nuccioreggie):

no you multiply oooooooohhhh its C @misty1212

OpenStudy (nuccioreggie):

@misty1212

OpenStudy (nuccioreggie):

@misty1212

OpenStudy (nuccioreggie):

@misty1212